# Diversity of disc viscosities can explain the period ratios of resonant and non-resonant systems of hot super-Earths and mini-Neptunes

### v1

@@ 4.2 Planetary growth
The growth of the planets is initially driven by pebble accretion (e.g. Johansen & Lacerda 2010; Ormel & Klahr 2010; Lambrechts & Johansen 2012). While pebble accretion is an efficient mechanism to grow planets, individual pebble-accreting planets are actually very inefficient in reducing the overall inward flux of pebbles (Lambrechts et al., 2014; Bitsch et al., 2019a), as long as they did not reach the pebble isolation mass. As the pebble flux is not reduced significantly by one growing planet, systems of multiple planets can efficiently form (Lambrechts et al., 2019; Bitsch et al., 2019a; Izidoro et al., 2021; Matsumura et al., 2021). In addition, the growth in the outer regions of the disc happens on longer time scales, so that the planets forming in the inner system have already reached the pebble isolation mass, diminishing the influence of more outer embryos on the growth of inner embryos (Bitsch et al., 2019a, 2020; Bitsch & Izidoro, 2023).
In our simulations, we used 30 planetary embryos, while the simulations of Izidoro et al. (2021) used initially 60 planetary embryos. Consequently, more planets can initially grow in the simulations of Izidoro et al. (2021), but they are also distributed in a wider radial space compared to our simulations. Furthermore, planets forming in the outer regions (e.g. exterior to 10 AU) will not migrate interior to 0.7 AU in low viscosity environments (see appendix A), and remain in the outer disc causing late instabilities of the inner systems. Consequently, we think that more embryos in the low viscosity simulations would not significantly change the outcome of our simulations in respect to the inner system structure (see also appendix D).
On the other hand, in the high viscosity simulations of Izidoro et al. (2021), planets migrate inwards faster and can thus reach the inner edge of the disc even if they originate from exterior of 10 AU. Consequently, more planetary embryos can migrate into the inner region of the disc. This can result in a squeeze of the formed planetary chains, where incoming planets can push already trapped planets into tighter resonances. In combination with the faster migration speed that can allow tighter resonances as well, the formed systems are tighter packed with more planets in the high viscosity simulations.
In appendix D, we show a comparison to the simulations of Bitsch & Izidoro (2023), who also performed simulations with 15 and 60 initial embryos. These simulations support the idea that the migration velocity, caused by changes in the disc’s viscosity, are responsible for the different chain configurations rather than the number of planetary embryos. In order to verify this result, we conducted simulations with only 2 planets in discs with different viscosities in appendix G, which confirm this result very clearly.
@@ 4.3 Formation of Kepler-223 and TOI-178
The Kepler-223 systems consists of four super-Earths, which are locked in a resonance chain, with 4:3, 3:2 and 4:3 resonances (Mills et al., 2016). The formation of a particular resonance chain requires detailed studies about the migration history of the planets. Hühn et al. (2021) executed an extensive parameter study within N-body simulations with prescribed migration rates to explain the migration history of Kepler-223. In particular, they varied the gas surface density value and the disc’s viscosity in order to find a parameter combination that allows the formation of that system.
Hühn et al. (2021) find that surface densities higher than the MMSN (Weidenschilling, 1977; Hayashi, 1981) (which are also higher than the gas surface densities used here) and -viscosities with are preferred to form the orbital configuration of Kepler-223. This result is in line with our low viscosity simulations that can reproduce a significant fraction of systems with 4:3 and 3:2 resonances. This result is encouraging considering the completely different set-up of the simulations of Hühn et al. (2021) and ours, indicating that lower viscosities could be preferred to produce systems in the 4:3 and 3:2 resonances.
The TOI-178 systems consists of 6 transiting planets following a 2:4:6:9:12 chain of Laplace resonances (Leleu et al., 2021). Our simulations indicate that these wide chains could be better explained by low viscosity environments, as the chains built in high viscosity environments can basically not produce planets in the 2:1 resonance configuration.
@@ 4.4 Consequences for inner and outer system structure
Our simulations show that the structure of the inner system is influenced by the outer regions (P200 days). While the systems formed in the low viscosity environments experience dramatic instabilities that can result in a large fraction of completely destroyed systems, the instabilities in the high viscosity systems mostly “just” break the chains, but leave the overall systems intact. We note that these instabilities do not necessarily eject the planets from the system. In fact, a large fraction of the planets is thrown into the host star, where they can influence the stellar atmospheric abundances (e.g. Liu et al. 2024). This difference is caused by the number and larger masses of planets in the outer disc regions (exterior to 1 AU, see Fig. 10), which trigger the violent instabilities in the systems formed in the low viscosity environments. This is in line with the simulations of Bitsch & Izidoro (2023), where outer giant planets destroy the inner systems, with the difference that the systems in Bitsch & Izidoro (2023) already become unstable during the gas-disc phase due to the larger impact of the giants compared to outer sub-Neptunes in the here presented simulations.
@@
This difference has two important implications for the occurrence rates of planets in the inner systems as well as in the outer systems. Even though all our simulations allow the formation of inner super-Earths during the gas disc stage, only around 50% of our final systems harbor inner planets after 100 Myr of integration. On the other hand, all our systems still harbor planets exterior to 1 AU after 100 Myr of integration. This result is broadly in line with the occurrence rate of sub-Neptunes exterior to the water ice line from microlensing surveys (e.g. Suzuki et al. 2016), which predict that these planets should be more common than hot inner super-Earths.
@@
We note that the presence of the outer systems (r¿1 AU) is influenced by our choice of the initial embryo distribution. If we were to include only a smaller number of embryos in our simulations (e.g. less than 8) within a few AU, all embryos would migrate into the inner disc regions and no outer system would exist. In this case, the instabilities would be less pronounced and the inner systems would survive, as they do in the high viscosity simulations. However, the lower viscosity would still determine the formed resonance chains, because the lower viscosity allow slower migration speeds, trapping the planets in wider configurations. This can be see in Fig. 4, where the period ratios of planets clearly show wider systems for the low viscosity simulations compared to the high viscosity simulations. On the other hand, if no outer embryos were to exist initially, it is unclear how the formation of sub-Neptunes exterior to the water ice line, as found in microlensing surveys (e.g. Suzuki et al. 2016), came to exist.
4.5 Comparison to other planet formation studies
@@
The formation of super-Earth and mini-Neptune systems has been studied by various groups in general (e.g. Ogihara et al. 2018; Emsenhuber et al. 2021; Ogihara et al. 2024; Batygin & Morbidelli 2023) or applied to specific systems (e.g. Coleman et al. 2019). These studies differ in their methods regarding accretion (pebbles or planetesimals), disc evolution (viscous or disc winds), or initial embryo formation (in ring like structures or distributed all over the disc), but they all include planetary migration, which is necessary for the build up of resonance chains. More noteably, most studies use the migration prescription by Paardekooper et al. (2011), as we do in our here presented work. Furthermore, all simulations include an inner edge, where inward migration can stop and planets can pile up (e.g. Masset et al. 2006; Flock et al. 2019).
@@
However, independently how the planets exactly form (pebbles or planetesimals) or in what configuration they form (in rings or wide spread), they have to migrate to the inner edge of the protoplanetary disc. We expect that migration speed, determined by the viscosity (as well as by the gas surface density and temperature, see appendix B) is the crucial component that will determine the resonance configuration that the planets will end up with. This is motivated by the fact that capture into resonances is determined by the relative migration velocity and not by the initial distance between the embryos.
@@
On the other hand, as larger viscosities allow the inward migration from planets from beyond the water ice line, a mixture in the composition of the systems are expected (e.g. Izidoro et al. 2022). At low viscosity, this might not be the case, unless the discs are very cold and the water ice line is initially close to 1 AU (e.g. Bitsch et al. 2019b) allowing planets to reach the inner disc even in low viscosity environments (see also appendix A). However, it is still under debate how much of water ice pebbles could be retained by growing planets, as the pebbles might evaporate in the hot atmosphere of planets, preventing the efficient accretion of water ice (Johansen et al., 2021; Müller et al., 2024).
@@ 5 Summary
We have conducted simulations similar to Izidoro et al. (2021), but using a low viscosity environment. This low viscosity environment allows planets of a few Earth masses to already open partial gaps, reducing their inward migration rate. Consequently, the planetary chains that form at the inner edge of the disc are in wider configurations compared to planets formed in a high viscosity environment. This has consequences for the stability of the systems and for their observabilities. Our findings are as follows:
@@
The lower viscosity causes slower migration speeds and consequently planets remain in wider configurations after the gas disc phase, where planets are mostly trapped in the 4:3, 3:2 and 2:1 resonance. The higher viscosity simulations, on the other hand, do not populate the 2:1 resonance directly after the gas disc phase and only marginally populate the 3:2 resonance. Most planets are in the 4:3 or even tighter configurations.
@@
The overall period ratios of the low viscosity simulations, after instabilities, remain too narrow to explain the observed period ratios. In contrast, the high viscosity simulations give a overall better match to the overall period ratio distribution (Izidoro et al., 2021).
@@
The observed multi-planet chains in the low viscosity environment, on the other hand, match very nicely with the observations. In particular, our low viscosity simulations can reproduce the large number of planets in the 3:2 and 2:1 configurations. This result is also not affected by outer giant planets that can destabilize the inner systems (Bitsch & Izidoro, 2023), indicating that low viscosity migration might be the cause of systems with these wider resonances, compared to the 5:4 (or tighter) resonances, which should form in discs with high viscosities.
@@
Obviously a large variety of parameters influence the outcome of planet formation simulations. Here we have identified that different migration rates (caused by different viscosities) have consequences on the chains that are formed in these simulations. As the low viscosity simulations match the chains better compared to the high viscosity simulations and as the high viscosity simulations match the overall observed period ratios better compared to the low viscosity simulations, we suggest that the majority of the resonance chains might be formed at low viscosity, most likely without nearby external perturbers, as these tend to destroy the resonance chains effectively. In contrast, the majority of the systems in non-resonant configuration might be formed in high viscosity environments with a small number of external perturbers that allow a breaking of chains rather than a destruction of chains. This interpretation indicates that a variety of viscosities (and outer companions) is probably needed to explain the different observed systems. Furthermore, this result is independent on the exact formation mechanism of planets (e.g. pebbles or planetesimals) as well as on the initial embryo configuration, as long as the embryos can migrate and form resonance chains at the inner edge of the disc.
@@ Acknowledgements.
A. I. acknowledges NASA grant 80NSSC18K0828 to Rajdeep Dasgupta, during preparation and submission of the work. A. I. also acknowledges support from the Welch Foundation grant No. C-2035-20200401. We thank the referee for their comments that helped us to improve our manuscript.References
@@
- Batygin & Morbidelli (2023) Batygin, K. & Morbidelli, A. 2023, Nature Astronomy
@@ Appendix A Migration of single planets
We show in Fig. 8 the semi-major axis evolution of 10 Earth mass planets as a function of time. The planets start at either 1 or 10 AU and are embedded in discs with different viscosities. The planets at low viscosity migrate the slowest, while planets at higher viscosity migrate fastest, due to partial gap opening at low viscosity. Consequently, we expect the planets in the high viscosity regime to result in tighter configurations due to the larger mutual migration velocities (see appendix G).
@@ Appendix D Influence of embryo number and disc viscosity
On the other hand, if only 15 initial embryos are used, the situation is different. Here we see a pile-up mostly in the 2:1 resonance configuration, rather than the 3:2 and 4:3. This is caused by the wide separation of the planetary embryos, where the embryos are spaced initially around 1.0 AU apart (Bitsch & Izidoro 2023). Consequently only 1-2 planets reach the inner edge of the disc, where the outer planet already transitioned into a gas giant, which reduced their migration speed significantly so that the 2:1 configuration can not be overcome. Furthermore, the large initial separation prevents the a 2nd planet to reach the inner edge, so that in most cases no resonant system is formed.
In Fig. 12 we show the period ratios of our simulations with , as discussed in the main part of the paper, as well as from comparison simulations using the same disc and planet parameters, but . Both sets of simulations feature 30 initial embryos. We show the periods at the end of the gas disc lifetime of 3 Myr. Clearly, the planetary systems formed in the high viscosity simulations result in tighter resonant chains compared to the systems formed in the low viscosity environments. In particular, the high viscosity simulations feature a larger fraction of planets in the 4:3 and 5:4 (and tighter) resonance configurations, while the low viscosity simulations produce many more planets in the 3:2 and 2:1 resonance configurations. We note that the exact period configurations are subject to change after the instability phase, but instabilities break the resonances and do not populate the resonance configurations, indicating that overall trend that higher viscosity systems produce resonant chains in tighter configurations compared to low viscosity simulations will continue to hold.

### v2

@@ 4.2 Planetary growth
The growth of the planets is initially driven by pebble accretion (e.g. Johansen & Lacerda 2010; Ormel & Klahr 2010; Lambrechts & Johansen 2012). While pebble accretion is an efficient mechanism of planet growth, individual pebble-accreting planets are actually very inefficient in reducing the overall inward flux of pebbles (Lambrechts et al., 2014; Bitsch et al., 2019a), as long as they do not reach the pebble isolation mass. As the pebble flux is not reduced significantly by one growing planet, systems of multiple planets can efficiently form (Lambrechts et al., 2019; Bitsch et al., 2019a; Izidoro et al., 2021; Matsumura et al., 2021). In addition, the growth in the outer regions of the disc happens on longer timescales, meaning that the planets forming in the inner system reach the pebble isolation mass more quickly than those forming in the outer regions, diminishing the influence of more outer embryos on the growth of inner embryos (Bitsch et al., 2019a, 2020; Bitsch & Izidoro, 2023).
In our simulations, we used 30 planetary embryos, while the simulations of Izidoro et al. (2021) used initially 60 planetary embryos. Consequently, more planets can initially grow in the simulations of Izidoro et al. (2021), but they are also distributed in a wider radial space compared to our simulations. Furthermore, planets forming in the outer regions (e.g. exterior to 10 AU) will not migrate interior to 0.7 AU in low-viscosity environments (see Appendix A), and remain in the outer disc, causing late instabilities of the inner systems. Consequently, we think that a greater number of embryos in the low-viscosity simulations would not significantly change the outcome of our simulations with respect to the inner system structure (see also Appendix D).
On the other hand, in the high-viscosity simulations of Izidoro et al. (2021), planets migrate inwards faster and can thus reach the inner edge of the disc even if they originate from beyond 10 AU. Consequently, more planetary embryos can migrate into the inner region of the disc. This can result in a squeeze of the formed planetary chains, where incoming planets can push already trapped planets into tighter resonances. In combination with the faster migration speed, which can allow tighter resonances as well, the formed systems are more tightly packed, with more planets in the high-viscosity simulations.
In Appendix D, we show a comparison to the simulations of Bitsch & Izidoro (2023), who also performed simulations with 15 and 60 initial embryos. These simulations support the idea that the migration velocity —set by disc viscosity— is responsible for the different chain configurations rather than the number of planetary embryos. In order to verify this result, we conducted simulations with only two planets in discs with different viscosities in Appendix G, confirming this phenomenon in simulations very clearly.
@@ 4.3 Formation of Kepler-223 and TOI-178
The Kepler-223 system consists of four super-Earths locked in a resonance chain with 4:3, 3:2, and 4:3 resonances (Mills et al., 2016). Understanding the formation of a particular resonance chain requires detailed studies of the migration history of the planets. Hühn et al. (2021) carried out an extensive parameter study using N-body simulations with prescribed migration rates to explain the migration history of Kepler-223. In particular, these authors varied the gas surface density value and the disc viscosity in order to find a parameter combination that allows the formation of that system.
Hühn et al. (2021) find that surface densities higher than the Minimum Mass Solar Nebular, or MMSN (Weidenschilling, 1977; Hayashi, 1981) (which are also higher than the gas surface densities used here) and -viscosities with favour the formation of the orbital configuration of Kepler-223. This result is in line with our low-viscosity simulations, which can reproduce a significant fraction of systems with 4:3 and 3:2 resonances. This result is encouraging considering the completely different setup of the simulations of Hühn et al. (2021) and ours, indicating that lower viscosities could favour the production of systems in the 4:3 and 3:2 resonances.
The TOI-178 system consists of six transiting planets following a 2:4:6:9:12 chain of Laplace resonances (Leleu et al., 2021). Our simulations indicate that such wide chains could be better explained by low-viscosity environments, as the chains built in high-viscosity environments do not produce planets in the 2:1 resonance configuration.
@@ 4.4 Consequences for inner and outer system structure
Our simulations show that the structure of the inner system is influenced by the outer regions (P200 days). While the systems that form in the low-viscosity environments experience dramatic instabilities that can result in a large fraction of completely destroyed systems, the instabilities in the high-viscosity systems mostly simply break the chains but leave the overall systems intact. We note that these instabilities do not necessarily eject the planets from the system; indeed, a large fraction of the planets are thrown into the host star, where they can influence the stellar atmospheric abundances (e.g. Liu et al. 2024). This difference is caused by the greater number and larger masses of planets in the outer disc regions (exterior to 1 AU; see Fig. 10), which trigger the violent instabilities in the systems formed in the low-viscosity environments. This is in line with the simulations of Bitsch & Izidoro (2023), where outer giant planets destroy the inner systems, with the difference that the systems in Bitsch & Izidoro (2023) already become unstable during the gas-disc phase due to the greater impact of the giants compared to that of the outer sub-Neptunes in the simulations presented here.
@@
This difference has two important implications for the occurrence rates of planets in the inner systems as well as in the outer systems. Even though all our simulations allow the formation of inner super-Earths during the gas-disc stage, only around 50% of our final systems harbour inner planets after 100 Myr of integration. On the other hand, all our systems still harbour planets exterior to 1 AU after 100 Myr of integration. This result is broadly in line with the occurrence rate of sub-Neptunes exterior to the water-ice line from microlensing surveys (e.g. Suzuki et al. 2016), which predict that these planets should be more common than hot inner super-Earths.
@@
We note that the presence of the outer systems (r¿1 AU) is influenced by our choice of the initial embryo distribution. If we were to include only a small number of embryos in our simulations (e.g. less than 8) within a few astronomical units (AU), all embryos would migrate into the inner disc regions and no outer system would exist. In this case, the instabilities would be less pronounced and the inner systems would survive, as they do in the high-viscosity simulations. However, the lower viscosity would still determine the formed resonance chains, because the lower viscosity allows slower migration speeds, trapping the planets in wider configurations. This can be seen in Fig. 4, where the period ratios of planets clearly show wider systems for the low-viscosity simulations compared to the high-viscosity simulations. On the other hand, if no outer embryos were to exist initially, it is unclear how the sub-Neptunes exterior to the water-ice line, as found in microlensing surveys (e.g. Suzuki et al. 2016), would come to exist.
4.5 Comparison to other planet-formation studies
@@
The formation of super-Earth and mini-Neptune systems has been studied by various groups in general (e.g. Ogihara et al. 2018; Emsenhuber et al. 2021; Ogihara et al. 2024; Batygin & Morbidelli 2023) and specific systems have also been examined (e.g. Coleman et al. 2019). These studies differ in their methods regarding accretion (pebbles or planetesimals), disc evolution (viscous or disc winds), and initial embryo formation (in ring-like structures or distributed all over the disc), but they all include planetary migration, which is necessary for the build up of resonance chains. More notably, most studies use the migration prescription by Paardekooper et al. (2011), as we do in the work presented here. Furthermore, all simulations include an inner edge, where inward migration can stop and planets can pile up (e.g. Masset et al. 2006; Flock et al. 2019).
@@
However, independently of how the planets form (pebbles or planetesimals) or the configuration in which they form (in rings or widely spread), they have to migrate to the inner edge of the protoplanetary disc. We believe migration speed, which is determined by the viscosity of the gas disc (as well as by the gas surface density and temperature; see Appendix B), to be the crucial component that determines the eventual resonance configuration of the planets. This belief is motivated by the fact that capture into resonances is determined by the relative migration velocity and not by the initial distance between the embryos.
@@
On the other hand, as larger viscosities allow the inward migration of planets from beyond the water-ice line, a mixture in the composition of the systems is expected (e.g. Izidoro et al. 2022). At low viscosity, this may not be the case, unless the discs are very cold and the water-ice line is initially close to 1 AU (e.g. Bitsch et al. 2019b), allowing planets to reach the inner disc even in low-viscosity environments (see also Appendix A). However, how much of the water ice in pebbles can be retained by growing planets is still debated, as the pebbles might evaporate in the hot atmosphere of the planets, preventing the efficient accretion of water ice (Johansen et al., 2021; Müller et al., 2024).
@@ 5 Summary
We conducted simulations similar to those of Izidoro et al. (2021), but using a low-viscosity environment. This low-viscosity environment allows planets of a few Earth masses to already open partial gaps, reducing their rate of inward migration. Consequently, the planetary chains that form at the inner edge of the disc are in wider configurations compared to planets that form in a high-viscosity environment. This has consequences for the stability of the systems and for their observability. Our findings are as follows:
@@
The lower viscosity causes slower migration speeds and consequently planets remain in wider configurations after the gas-disc phase, where planets are mostly trapped in the 4:3, 3:2, and 2:1 resonances. The higher-viscosity simulations, on the other hand, do not populate the 2:1 resonance directly after the gas-disc phase and only marginally populate the 3:2 resonance. Most planets are in the 4:3 resonance or exhibit an even tighter configuration.
@@
The overall period ratios of the low-viscosity simulations, after instabilities, remain too narrow to explain the observed period ratios. In contrast, the high-viscosity simulations, in general, provide a better match to the overall period ratio distribution (Izidoro et al., 2021).
@@
On the other hand, the simulated multi-planet chains in the low-viscosity environment very closely match observations. In particular, our low-viscosity simulations can reproduce the large number of planets in the 3:2 and 2:1 configurations. This result is also not affected by outer giant planets, which can destabilise the inner systems (Bitsch & Izidoro, 2023), indicating that low-viscosity migration might be the major driver of wider resonances than the 5:4 (or tighter) resonances, which should therefore be more likely to form in discs with high viscosities.
@@
Clearly, a large variety of parameters influence the outcome of planet-formation simulations. Here we have identified that different migration rates (caused by different viscosities) have consequences on the chains that are formed in these simulations. As the low-viscosity simulations better match the chains than the high-viscosity simulations, and as the high-viscosity simulations better match the overall observed period ratios than the low-viscosity simulations, we suggest that the majority of the resonance chains are most probably formed at low viscosity, most likely without nearby external perturbers, as these tend to destroy the resonance chains effectively. In contrast, the majority of the systems in non-resonant configurations are likely to have formed in high-viscosity environments with a small number of external perturbers that allow a breaking of the chains rather than their total destruction. This interpretation indicates that a variety of viscosities (and outer companions) is probably needed to explain the different observed systems. Furthermore, this result is independent of the exact formation mechanism of the planets (e.g. pebbles or planetesimals), and of the initial embryo configuration, as long as the embryos can migrate and form resonance chains at the inner edge of the disc.
@@ Acknowledgements.
We thank the referee for their comments that helped us to improve our manuscript.References
@@
- Batygin & Morbidelli (2023) Batygin, K. & Morbidelli, A. 2023, Nature Astronomy, 7, 330
@@ Appendix A Migration of single planets
We show in Fig. 8 the semi-major axis evolution of 10 Earth mass planets as a function of time. The planets start at either 1 or 10 AU and are embedded in discs with different viscosities. The planets at low viscosity migrate the slowest, while planets at higher viscosity migrate fastest, due to partial gap opening at low viscosity. Consequently, we expect the planets in the high-viscosity regime to result in tighter configurations due to the larger mutual migration velocities (see Appendix G).
@@ Appendix D Influence of embryo number and disc viscosity
On the other hand, if only 15 initial embryos are used, the situation is different. Here we see a pile-up mostly in the 2:1 resonance configuration, rather than the 3:2 and 4:3. This is caused by the wide separation of the planetary embryos, where the embryos are spaced initially around 1.0 AU apart (Bitsch & Izidoro 2023). Consequently only 1-2 planets reach the inner edge of the disc, where the outer planet already transitioned into a gas giant, which reduced their migration speed significantly so that the 2:1 configuration cannot be overcome. Furthermore, the large initial separation prevents the second planet from reaching the inner edge such that in most cases no resonant system is formed.
In Fig. 12 we show the period ratios of our simulations with , as discussed in the main part of the paper, as well as from comparison simulations using the same disc and planet parameters, but . Both sets of simulations feature 30 initial embryos. We show the periods at the end of the gas disc lifetime of 3 Myr. Clearly, the planetary systems formed in the high-viscosity simulations result in tighter resonant chains compared to the systems formed in the low-viscosity environments. In particular, the high-viscosity simulations feature a larger fraction of planets in the 4:3 and 5:4 (and tighter) resonance configurations, while the low-viscosity simulations produce many more planets in the 3:2 and 2:1 resonance configurations. We note that the exact period configurations are subject to change after the instability phase, but instabilities break the resonances and do not populate the resonance configurations, indicating that overall trend that higher viscosity systems produce resonant chains in tighter configurations compared to low-viscosity simulations will continue to hold.